OpenStudy (anonymous):

Find the product of the complex numbers. Express your answer in trigonometric form. z1 = 5(cos(25) + isin(25)) z2 = 2(cos(80) + isin(80)) A. z1 = 3(cos(-55) + isin(-55)) B. z1 = 10(cos(105) + isin(105)) C. z1 = 7(cos(105) + isin(105)) D. z1 = 10(cos(305) + isin(305))

OpenStudy (tkhunny):

#1 It makes no sense to have "z1" in all the answer, unless we're not supposed to notice that nothing matches the original definition. #2 How does DeMoivre work? If z1 were equal to z2, what would happen?

OpenStudy (tkhunny):

@eashy Please, never do that again. Read the code of conduct and consider compliance.

OpenStudy (anonymous):

?

OpenStudy (tkhunny):

From the Code of Conduct "Give Help, Not Answers I will encourage and guide those needing help, and not just give them an answer"

OpenStudy (anonymous):

ok. use the transformation e^ix=cos(x)+i sin(x) to convert multiplication into addition (cos(25) + isin(25))*(cos(80) + isin(80))=e^25i*e^80i=e^105i=cos(105) + isin(105)